# Nonstabilizerness Estimation using Graph Neural Networks

### v1

@@ 6.3 Stabilizer Rényi Entropy Estimation
6.4 The Importance of the Graph Representation
Given the improved performance of the GNN compared to the SVR model, it is important to investigate the source of this advantage. We identify two main differences between the GNN and SVR approaches. First, the GNN employs a graph-based representation combined with Transformer Convolutional layers, which aggregate local information from neighboring nodes. Second, the GNN leverages a deep learning architecture to process the global feature set, which is the same set used in the SVR model [lipardi2025study].
To examine the importance of the graph representation, we perform experiments by isolating the fully-connected neural network from the upper part of the model shown in Figure 5. This is a common process in artificial intelligence, known as ablation, where an individual component is obscured in order to evaluate its contributions to the whole complex models [cohen1988evaluation, newell1975tutorial]. In this setup, we disabled the Transformers Convolutional responsible for processing the graph-structured data. Consequently, we train only the fully connected neural network (NNs) on the global feature set. The dotted bars in Figure 12 and Figure 13 show the results obtained in this experimental setup for the extrapolation on qubit number and gate count, respectively. Table 3 compares directly the extrapolation performances of GNN an NN with respect to the SVR model [lipardi2025study].
6.5 Quantum Noise Effect on Stabilizer Rényi Entropy Estimation
One of the main motivations for adopting the Stabilizer Rényi Entropy (SRE) as a measure of nonstabilizerness in this work is its suitability for experiments on real quantum devices [oliviero2022measuring, niroula2024phase]. For this reason, in this section we assess the impact of quantum noise on the predictions of our GNN models. The RQC Noisy and TIM Noisy datasets have been generated specifically for this task. In these datasets, all quantum circuits are labeled with their $M_{2}$ values obtained by simulating quantum noise using the IBM Fake Oslo backend.
The experiments in this setup follow exactly the same procedure as those described in Section 6.3. Figure 15 reports the mean squared error (MSE) values for the extrapolation experiments with respect to the qubit number (Figure 15(a)) and the gate count (Figure 15(b)). The SVR model was trained following the procedure described in [lipardi2025study], including a grid search over the hyperparameters.
(a) Extrapolation on qubit number
(b) Extrapolation on circuit depth
We note that the noisy RQC and TIM datasets are smaller in size compared to their respective noiseless counterparts, given the high computational time required for their simulation. For this reason, a comparison of the results between the model predictions obtained in the noisy and noiseless scenarios has to be interpreted with caution, as part of the performance gap may be attributed to the reduced training set size rather than the effect of noise itself.
7 Conclusions
This article proposed a Graph Neural Network (GNN) approach for addressing the nonstabilizerness estimation problem in the supervised learning setting through three formulations of increasing difficulty. The first is the stabilizer state classification, which is defined independently of any specific nonstabilizerness measure. This formulation serves primarily as a benchmark to establish a consistent data generation procedure and to perform an initial assessment of the GNN model on an easier problem. The second is the SRE-based classification defined on the stabilizer Rényi entropy (SRE). This formulation represents a more challenging and physically motivated task, as it directly connects to identifying stabilizer states based on their SRE measured on a real quantum device. Finally, the SRE estimation, is the more general regression task, which aims to predict the SRE of quantum circuits as a quantitative measure of nonstabilizerness [lipardi2025quantum].
In the first two classification tasks, the GNN model is trained on product-state quantum circuits. Compared to previous CNN-based approaches [mello2025retrieving], the proposed GNN model achieves more robust generalization performance on circuits evolved under Clifford operations. Since the CNOT is a Clifford gate, Clifford-evolved circuits allow to test the generalization capabilities to entangled states. Our proposed data generation procedure includes not only stabilizer and highly magic states but also quantum states with low SRE values. This diversity is a fundamental feature of the dataset, as it prevents the problem from being defined on states that are easily separable. Additionally, we extended the analysis to circuits with different structure and high number of CNOT gates, as well as on circuits with higher number of qubits, where the GNN also exhibits robust performances. For the regression task, our GNN model provides more accurate SRE predictions than the Support Vector Regressor [lipardi2025study], especially in terms of generalization to larger, previously unseen quantum circuits. By ablation of the graph representation and the GNN components that process the graph data, we show that the graph representation improves the model predictions. The impact of the graph representation is significantly relevant on random quantum circuit where circuit structures are quite different. Moreover, the experimental results show that the proposed GNN is a promising approach to predict SRE values measured on a given hardware platform. This is because the graph-based framework effectively accommodates the backend information of quantum devices. However, more extensive experiments involving tests on different quantum devices and noise models are necessary to thoroughly assess the capability of our GNN to predict the noisy SRE measurements on real quantum devices. To facilitate benchmarking and support advancements in this field, we release a comprehensive dataset of quantum circuits specifically designed for the nonstabilizerness estimation problem in those three formulations.
The performance analysis on the experimental demonstrate that by leveraging the graph-based representation of quantum circuits, the proposed GNN model achieves accurate nonstabilizerness estimation with robust generalization capabilities to larger quantum circuits, in terms of both number of quantum circuits and gate count, and to higher entanglement. We conclude that the proposed GNN model offers an effective and scalable framework for estimating nonstabilizerness across both noiseless and noisy regimes.
Future Directions
We identify three main future research directions. The first regards the design of different features that can improve the graph-based quantum circuit representation for our GNN model. The integration of expectation values of local observables, which can be calculated efficiently through the classical shadows protocol [huang2020predicting], can be an interesting direction as suggested by previous theoretical [mello2025retrieving] and experimental results [lipardi2025study].
The third future direction is the integration of the proposed GNN model within Quantum Architecture Search (QAS) frameworks, to enable the development of SRE-informed QAS techniques [du2022quantum, zhang2022differentiable]. A particularly suitable framework to embed SRE estimation in quantum architecture search is Monte Carlo Tree Search [lipardi2025quantum], where the information regarding the nonstabilizerness can be used in the selection or expansion steps to push the search towards quantum circuits difficult to simulate on a classical device. This direction might be significantly beneficial in the quest for achieving quantum advantage with variational quantum algorithms (VQAs) [cerezo2021variational], as it would allow the design of quantum circuits that optimize both the quality of the solution to a target problem and their complexity in terms of classical simulation. Moreover, as evaluating objective functions in QAS requires typically significant computational effort, the proposed GNN can be used as a surrogate model to approximate these objectives. Such integration would enhance QAS efficiency and aligns with the pretraining framework described in the second future direction.
Data Availability
The data and code to reproduce the experiments described in the article are publicly available at Zenodo [lipardi2025dataset] and GitHub: https://github.com/VincenzoLipardi/GNN_SRE.
Acknowledgments
The authors acknowledge the financial support of the Quantum Software Consortium, which facilitated the development of this work and strengthened collaborations through a visitor travel grant.
Conflict of Interests
The authors declare no conflict of interests.

### v2

@@ B.2 Performance Analysis for the Stabilizer Rényi Entropy Estimation
B.3 Extended Analysis to the Extrapolation in Qubits
In this section, we extend our experiments to evaluate the generalization performance of the GNN on circuits with higher qubit counts, beyond those tested in Section 6.3 and in the previous work [32]. The limitations on the number of circuits arise from the computational time required to compute $M_{2}$ for completely unstructured circuits in the RQC dataset, which increases exponentially with the number of qubits. Nevertheless, we expanded the RQC dataset to include circuits with up to $6$ qubits, adding $200$ additional circuits equally divided between $7$- and $8$-qubit circuits.
The GNN described in Section 6.3 is trained on circuits from the RQC dataset containing up to $5$ qubits. The extrapolation MSE reported in Figure 7 is computed on the remaining $6$-qubit circuits, which are $50000$. Figure 16 recalls the extrapolation MSE achieved on the $6$-qubit circuits and extends the analysis to the newly added $7$- and $8$-qubit circuits.
We observe that the extrapolation MSE achieved on the few additional circuits is consistent with that obtained on the more statistically significant set of $6$-qubit circuits. The reduced extrapolation MSE observed for the $7$-qubit circuits can likely be attributed to a favorable random generation of circuits. In general, the extrapolation MSE increases for circuits with a number of qubits farther from the training set, although it remains reasonably stable overall. This observation also aligns with the experiments on generalization performance with respect to the number of qubits in the classification setting.
Figure 16: Extrapolation Mean Squared Error (MSE) in the number of qubits for the GNN.
Appendix C. Generalization to Structured Quantum Circuits
In this section, we study the generalization capabilities of the proposed GNN approach when trained on the unstructured RQC dataset and tested on two different structured classes of quantum circuits. The first is derived by the one-dimensional TIM model and the second one-dimensional Heisenberg model. Similarly to the TIM model, which we extensively discussed in Appendix A.3 Dataset for Stabilizer Rényi Entropy Estimation, in the Heisenberg model, each quantum circuit represents a discrete-time approximation of the continuous-time evolution generated by the Heisenberg Hamiltonian. For a one-dimensional chain of $n$ qubits with nearest-neighbor interactions, the Hamiltonian is given by:
$H_{Heis}=\sum_{i=1}^{n-1}J_{X}X_{i}X_{j}+J_{Y}Y_{i}Y_{j}+J_{Z}Z_{i}Z_{j}$ (7)
where $J_{X}$, $J_{Y}$, and $J_{Z}$ are real coupling constants.
The corresponding time-evolution operator $U(t)=e^{-itH_{\mathrm{Heis}}}$ is approximated using a first-order Trotter–Suzuki decomposition [38]:
$U(t)\approx\left[\prod_{i=1}^{n-1}e^{-i\theta_{X}X_{i}X_{i+1}}e^{-i\theta_{Y}Y_{i}Y_{i+1}}e^{-i\theta_{Z}Z_{i}Z_{i+1}}\right]^{T},$ (8)
where $\theta_{\alpha}=J_{\alpha}\Delta t$ for $\alpha\in\{X,Y,Z\}$. Each two-qubit interaction term is implemented using elementary quantum gates. The $ZZ$ interaction $e^{-i\theta_{Z}Z_{i}Z_{i+1}}$ is realized by the gate sequence $\texttt{CNOT}_{i,i+1}$$\texttt{RZ}(2\theta_{Z})_{i+1}$$\texttt{CNOT}_{i,i+1}$. The $XX$ interaction $e^{-i\theta_{X}X_{i}X_{i+1}}$ is obtained by conjugating the $ZZ$ interaction with Hadamard gates on both qubits, while the $YY$ interaction $e^{-i\theta_{Y}Y_{i}Y_{i+1}}$ is implemented by conjugating the $ZZ$ interaction with $R_{X}(\pi/2)$ and $R_{X}(-\pi/2)$ rotations. The circuits span various values of the angle parameters $\theta_{\alpha}$, with the number of Trotter steps $T$ ranging from one to five.
The experimental results indicate that when the GNN is trained on the unstructured RQC dataset, the predictions do not generalize well to structured circuits. However, by fine-tuning the model on a small number of circuits from the corresponding class of quantum circuits, the mean squared error (MSE) converges to values comparable to those reported in Section 6.3. Figure 17 illustrates the extrapolation MSE on both structured datasets under the configuration in which testing is performed on unseen numbers of qubits and gate counts. We observe that fine-tuning on as few as $50$ circuits leads to a substantial decrease in MSE, and after $500$ circuits, the MSE approaches the values reported in Figure 7 and Figure 8, where the training set includes $16000$ circuits.
Figure 17: Extrapolation MSE on two class of structured quantum circuits obtained by the GNN trained on the unstructured RQC dataset.
Appendix D. Benchmark against Analytic Estimations
In this section, we compare the $M_{2}$ estimates obtained by the proposed GNN to those derived from an analytical formula for the stabilizer linear entropy presented in [29].
The stabilizer linear entropy $M_{\mathrm{lin}}$ of a quantum state $\rho$ is defined as
$M_{\mathrm{lin}}(\rho):=1-d\left\|\Xi(\rho)\right\|_{2}^{2},$
[…]
$\mathbb{E}_{C_{k}}\!\left[M_{\mathrm{lin}}(U)\right]=1-\frac{1}{3+d}\left(4+(d-1)\,f(\theta)^{k}\right),$ (9)
where $\theta$ is the angle of the Phase gate introducing magic in the circuit and
$f(\theta)=\frac{7d^{2}-3d+d(d+3)\cos(4\theta)-8}{8(d^{2}-1)}$
satisfies $f(\theta)\leq 1$, with equality for Clifford gates. Figure 18(a) shows $f(\theta)$ for representative values of $d$.
(a) $f(\theta)$ function from Equation 9
(b) $M_{lin}$ value distribution
Figure 18: $k$-doped random quantum circuits generated to compare the GNN prediction with the analytic estimation.
To benchmark the GNN against this analytic prediction, we generate a benchmark dataset of $4$-qubit circuits. We first sampled 100 random Clifford base circuits (single-qubit Clifford rotations and CNOT gates). For each base circuit, we generate a family of doped variants by inserting $k$ non-Clifford phase operations, with $k\in\{1,\dots,50\}$. This results in $100$ circuits per $k$ and $5000$ circuits in total. Non-Clifford doping is implemented exclusively as $R_{z}(\theta)$ gates (equivalent to $P(\theta)$ up to a global phase), placed in random qubits and position. Each base circuit is assigned a fixed random non-Clifford angle $\theta$ (uniformly sampled in $0,2\pi)$), shared across all its $k$-variants to analyze the performance in terms of the doping depth. Each circuit was labeled with its $M_{2}$ value, and $M_{\mathrm{lin}}$ was obtained from $M_{\mathrm{lin}}=1-e^{-M_{2}}$. The resulting $M_{\mathrm{lin}}$ distribution as a function of the doping depth $k$ is shown in Figure 18(b).
On this set of $k$-doped circuits, we benchmark the GNN (trained on the general RQC dataset) against the analytic estimate derived from Eq. 9. Figure 19 reports the comparison in terms of the mean squared error (MSE) with respect to the true $M_{\mathrm{lin}}$. The results indicate that the GNN achieves lower MSE than the analytic prediction. As shown in Fig. 19(b), the MSE of the GNN estimate (red line) consistently lies below that of the analytic estimate (blue line). The improvement can be attributed to the more informative input representation exploited by the GNN. Specifically, the GNN operates on a graph-based representation encoding the full circuit structure, including the angles and positions of all gates. In contrast, the analytic formula depends only on the dimension of the Hilbert space $d=2^{n}$, the number of non-Clifford gates $k$, and the angle $\theta$ of the Phase gate that injects magic into the circuit. Figure 19(a) illustrates the performance in terms of MSE as a function of $\theta$. The dataset contains $100$ distinct values of $\theta$, for better readability, the figure reports the averaged trend. Although the GNN generally outperforms the analytic estimate, the latter achieves lower error near the maxima of the function $f(\theta)$ (corresponding to Clifford gates), where it locally provides more accurate predictions than the GNN.
(a) MSE vs $\theta$
(b) MSE vs $k$
Figure 19: Performance of the GNN model compared to the analytic estimation based on Equation 9.